OpenStudy (anonymous):

What is the slope of a line given with this condition : 3x-2y+2 = 0 ??

OpenStudy (hoblos):

can you find the slope-intercept form by yourself??

OpenStudy (anonymous):

3x-2y+2=0 -2y=-3x-2 y=-3/2x+1 slope = -3/2 intercept (0,1)

OpenStudy (anonymous):

@hoblos yeah ! i'm just checking if it's right because i really don't get it i try to solve but my answer is really not right even if i try and try..

OpenStudy (hoblos):

if you have more questions post them and try to solve and tell me what you get so that i correct it for you and help you understand it ;)

OpenStudy (anonymous):

Make y the subject of the formula.Afterwards equation of a line is y+mx+c .M will be your slope.